OpenStudy (anonymous):

Identify the center and radius of the circle (x-5)^2+(y+1)^2=4

OpenStudy (anonymous):

A. Center: (-5, 1), r = 4 B. Center: (5, -1), r = 2 C. Center: (5, -1), r = 4 D. Center: (-5, 1), r = 2

OpenStudy (anonymous):

It's B. (x-h)^2+(y-k)^2=r^2 (h,k) = center r^2 = radius

OpenStudy (anonymous):

Ah, I see! Thank you for the help! (:

OpenStudy (anonymous):

No problem. ^^